OpenStudy (anonymous):

Which of the following is equivalent to |5x – 1| = 20?5x – 1 = 20.... 5x – 1 = 20 or 5x – 1 = -20 5x – 1 = -20 none of the above

OpenStudy (anonymous):

is it 5x – 1 = 20 or 5x – 1 = -20

OpenStudy (cwrw238):

5x - 1 = 20 , 5x - 1 = -20

OpenStudy (anonymous):

thnks

OpenStudy (cwrw238):

yw